OpenStudy (anonymous):

Can someone check my answers? A cube is 15cm on each side. Calculate its volume in cm^3, m^3, and mm^3. 3375cm^3 0.00375m^3 337500mm^3

OpenStudy (anonymous):

Can you tell me how to do this one? A cube is 10 inches on each side. Calculate its volume in m^3, cm^3, and dm^3. I'm not sure how to convert in^3 to m^3?

OpenStudy (cuanchi):

hold down 3375 cm^3 0.003375 m^3 3375000 mm^3

OpenStudy (anonymous):

Thanks!

OpenStudy (cuanchi):

\[10 inches \times10 inches \times10 inches =1000 inches ^{3}\] you can use the conversion factor 1inch= 2.54 cm to the power of 3 to made all cubic \[1000inches ^{3}\times \left( \frac{ 2.54 cm }{ 1inch } \right)^{3}=16387.064 cm ^{3}\]

OpenStudy (anonymous):

You are awesome! Thanks for breaking it down for me!